OpenStudy (anonymous):

Will give medal for any help!There were 300 people at a play. Some paid $5.00 for reserved seats; others paid $3.00 for general seats. Total receipts were $1,080. How many tickets of each kind were sold?

OpenStudy (texaschic101):

r + g = 300 5r + 3g = 1080 these are the equations you will need. Can you solve it from here ? It can be easily solved with substitution.

OpenStudy (texaschic101):

r + g = 300 r = 300 - g now sub 300 - g in for r in the other equation and you will find g. Then sub that back into either of the equations to find r. Does that make sense to you ?